Circular No. 6754

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1997 T1 (UTSUNOMIYA) Total visual magnitude estimates: Oct. 6.75 UT, 10.5 (M. Plsek, Lele-

kovice, Czech Republic, 0.35-m reflector; 10′ tail in p.a. 110^{o}; 6.76, 10.3 (V. Znojil, Brno, Czech Republic, 25 × 100 binoculars; 3′ tail in p.a. 140^{o}); 7.02, 10.0 (J. E. Bortle, Stormville, NY, 0.41-m reflector); 7.83, 9.8 (M. Reszelski, Szamotuly, Poland, 20 × 60 binoculars).

COMET C/1997 T2 (SOHO) C. St. Cyr, Naval Research Laboratory, on behalf of the SOHO-LASCO

Consortium (cf. IAUC 6685), reports that D. Lewis, University of Birmingham, has discovered a comet in both C3 and C2 coronagraphic data. The peak brightness was perhaps mag ∼ 6, and there was evidence of a small tail. Measurements by Lewis, D. Biesecker and St. Cyr have been reduced by G. V. Williams and are published in detail on MPEC 1997-T05, together with a Kreutz-type orbital solution by the undersigned.

1997 UT  α_{2000}  δ_{2000}
Oct. 3.287  12^{h}12.^{m}8  − 5^{o}20′

COMET 1997 T3 Uri Carsenty and Andreas Nathues, DLR Institute of Planetary Ex-

ploration, Berlin, report their discovery of a cometary object in the course of the Uppsala-DLR Trojan Survey program carried out at the European Southern Observatory together with C.-I. Lagerkvist, S. Mottola and G. Hahn. The object was detected on an exposure with the 1-m Schmidt as a point source, although seven unfiltered CCD images obtained on three consecutive nights with the Bochum 0.6-m reflector clearly show a tail extending 15′′ west-southwestward from the point source, which is estimated at R ∼ 18.9. E. Braatz also assisted with the Oct. 7 observations.

1997 UT  α_{2000}  δ_{2000}
Oct. 5.07800  1^{h}15^{m}20.^{s}50  +3^{o}04′03′′.0
5.08532  1 15 20.28  +3 04 01.6
6.21191  1 14 46.74  +3 01 10.9
6.21799  1 14 46.55  +3 01 10.1
7.15794  1 14 18.59  +2 58 47.9
7.16691  1 14 18.32  +2 58 46.5
7.17177  1 14 18.17  +2 58 45.8
1997 October 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden